OpenStudy (anonymous):

evaluate log[(tan2)(tan3)....(tan88)]

OpenStudy (anonymous):

0

OpenStudy (anonymous):

how ?

OpenStudy (anonymous):

log(tan2cot2tan3cot3.............) =log(1)

OpenStudy (anonymous):

i am beginner in trigonometry and logarithm .. please explain it further

OpenStudy (anonymous):

tan(88)=cot(2) because \[\tan(90-\theta)=\cot \theta\]

OpenStudy (anonymous):

thanks and then........

OpenStudy (anonymous):

you can find such angles all throughout except 45 but tan 45=1

OpenStudy (anonymous):

thnx

OpenStudy (anonymous):

i got it thanks very much